This is from MYMATHLAB Book from school
Find an equation of the line containing the given pair of points. Express your answer in the form x=a, y=b, or y=mx+ b
(-5, -5) and (5,5)
What is an equation of the line?
y=
Simplify your answer

Answer by stanbon(48546)

Find an equation of the line containing the given pair of points. Express your answer in the form x=a, y=b, or y=mx+ b
(-5, -5) and (5,5)
What is an equation of the line?
--------------------
slope = (5--5)/(5--5) = 1
Form is y = mx + b ; you know x,y and m; solve for b:
5 = 1*5 +b
b = 0
----------
EQUATION:
y = x
============
Cheers,
Stan H.